What is the APY for an account with a nominal annual rate of 5% compounded quarterly?
Answer #1

APY is calculated using below formula:

APY = (1 + r/n )n – 1

Annual interest rate = r = 5% = 0.05

Number of compounding period = n =Quarterly = 4

APY = (1 + 0.05/4)4 -1

        = (1 + 0.0125)4 -1

       = (1.0125)4 -1

       = 1.0509453 – 1

       = 0.0509453

       = 5.095 % (Approx)

The APY is 5.095%


source: https://www.bankrate.com/glossary/a/apy-annual-percentage-yield/
answered by: Sam Patil